To compile the code, you need to first create a directory in your unix account to hold the src code, binaries, etc. I will assume that you do the following:

cd
mkdir ncgb.example

Now find the nw directory. It should be in /home/ncalgdev/nw. Move it into your directory.

mv nw ncgb.example

By the way, nw stands for ”noweb”. Now you need to get into the nw directory.

cd ncgb.example
cd nw

Depending on what type of NCGB you want (for UNIX? for PC? with or without Mathematica?), you will need to change the options target in the Makefile. First find the target as follows

vi Makefile
/options:

Now, you should see something like

options: euclidmma

Note the tab character after the colon. It MUST be there. You may change this name from euclidmma. The possibilities are: •euclidmma ( with mma and on unix) •euclid ( standalone and on unix) •eulermma ( with mma and ???) •euler ( standalone and ???) •pcmma ( with mma and on PC with visual C++) •pc ( standalone and on PC with visual C++)

Save the file ’Makefile’ and exit vi. Now make the ’all target

make all

Now, if you are compiling for unix, just sit back, relax and drink coffee. If you are using unix

If the compile is successful, then you should move nw out of the current directory.

mv nw ../

Celebrate with NCGB. If you using a PC

If you are compiling for the PC, then there is more work to be done. After you type ’make all’ various things will happen and (eventually) there will be lines on the screen which start with ’g++’. At that point, you should type Control-C. If you don’t, it will compile a few files for a while and g++ will stop.

You should move nw out of the current directory.

mv nw ../

You then need to move to your pc. PC — Using samba

Connect up to samba with the help of a local expert. If you are UCSD, Dell knows how to do it. Click on windows until you get to a directory “\\MyMachine\myaccount\ncgb.examplebackslashnc\ncgb\compile”. Continue at Once in directory. PC — Not using samba

You need to use ftp. I will assume that you have WS FTP. Set up a directory on your pc. Start WS FTP. Mave the pc directories point to the directory which you just set up. Mave the unix directory point to your home directory.

Select ncalg.example.

Now hit the arrow from unix to the PC (the arrow in the middle) and answer yes that you want to move the directory. Wait for a few minutes.

Continue at Once in directory. Once in directory

Look for NCGB.dsw or NCGB.dsp. Click on one of them (if you pick the wrong one, VC++ will complain and ask to use the other. Just say ”ok”). Now click on the compile button. It is likely that your PC will freeze from 1 to 5 minutes and not appear to do anything. It will then start compiling. It takes less than an hour. Drink coffee or prove theorems.